OpenStudy (anonymous):

Solve (32/81)=(4/9)m

OpenStudy (anonymous):

m = 32/81 / 4/9 = 32/81 * 9/4 = 8 /9

OpenStudy (anonymous):

(9)* (32/81)=(4/9)m*(9) (32/9)/4=(4m)/4 8/9=m